OpenStudy (anonymous):

Find the slope of the graph of the function y=5x/x-2 at (3,15). Then find an equation for the line tangent to the graph at that point.

OpenStudy (anonymous):

@clavoie slope of the graph can be determined if we differentiate y w.r.t x

OpenStudy (anonymous):

wouldn't it be 5?

OpenStudy (anonymous):

@clavioe have u applied the quotient rule?

OpenStudy (anonymous):

I don't know...I'm taking this class online and I'm so lost right now

OpenStudy (anonymous):

If the function f(x) = g(x)/h(x) where h(x) is not equal to zero then according to quotient rule, f '(x) = h(x) g'(x) - g(x) h'(x) / {h(x)}^2

OpenStudy (anonymous):

try to apply this rule and u will finally get the slope

OpenStudy (anonymous):

I have no idea what that is or how to do that. We were given the equation \[\lim f(x _{0}+h)-f(x _{0})/h\] to solve these problems

OpenStudy (anonymous):

yeah it is first principle it will make the solution lil bigger can u help @gerryliyana

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you please just show me the steps on how to solve this?

OpenStudy (anonymous):

lets do this ques can u write f(x+h) ?

OpenStudy (anonymous):

f(3+h)= 5(3+h)/3+h-2

OpenStudy (anonymous):

cool now write f(x+h) - f(x) ?

OpenStudy (anonymous):

(5(3+h)/3+h-2)-(5(3)/3-2)

OpenStudy (anonymous):

can u simplify the above expression?

OpenStudy (anonymous):

(15+5h/1+h)-15

OpenStudy (anonymous):

take the LCM now and what will u get?

OpenStudy (anonymous):

I don't know, this is where I get stuck

OpenStudy (anonymous):

f(3+h)-f(3) = {15 + 5h - 15(1+h) } / (1+h) = {15 + 5h - 15 - 15h } / (1+h) = -10h/ (1+h)

OpenStudy (anonymous):

now f ' (3) = Lim f(3+h)- f(3)/h hope u will get the slope now

OpenStudy (anonymous):

So the slope is -10?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

in order to determine the equation , use point slope form according to point slope form y - y1 = m (x-x1) where m = slope of the line and (x1, y1) is any point on the line